what is the surface area of a cylinder if the radius is 2cm and the height is 5mm

Answer by Alan3354(69443)

2 cm = 20 mm
Lateral area = 2*pi*r*h = 200pi sq mm
Area of ends = pi*r^2 = 400pi sq mm each